A line passes through these points: (1,6) and (9, -4).
Could you please write an equation for the line in slope-intercept form and show me how step-by-step?
Thank you :)

Found 2 solutions by sudhanshu_kmr, stanbon:
Answer by sudhanshu_kmr(1152)

equation...
(y-6) = (6+4)/(1-9) (x-1)
=> y-6 = -5/4 (x-1)
=> 4y - 24 = -5x + 5
=> 4y = -5x + 29
=> y = -5/24 x + 29/4 ( in slope intercept form)

slope = -5/24 , y-intercept = 29/4

Answer by stanbon(75887)

A line passes through these points: (1,6) and (9, -4).
Could you please write an equation for the line in slope-intercept form and show me how step-by-step?
-------------------------
slope = (-4-6)/(9-1) = -10/8 = -5/4
------------------
Form: y = mx+b
You have y = 6 when x = 1 and you have m=-5/4
-------------------
6 = (-5/4)*1 + b
b = 6+(5/4)
---
b = 29/4
-------
Equation:
y = (-5/4)x+(29/4)
=====================
Cheers,
Stan H.
=============